# NGO Activism:

### v7

@@ Abstract
We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs). Using data on 2,500 U.S. campaigns, we find that NGOs often campaign on the AGM date, which increases media exposure and future shareholder proposals but does not affect current AGM votes or consumers. As NGOs gain reputational capital, they shift to earlier pre-AGM campaigns to support related shareholder proposals and directly improve corporate behavior. The evidence underscores a strategic trade-off between visibility and influence, showing how NGO tactics evolve with NGOs’ credibility and the changing expectations of socially conscious stakeholders about NGO activities.
@@ 1 Introduction
Non-governmental organizations (NGOs) are increasingly prominent actors in shaping firms’ environmental and social policies (Baron, 2001; Egorov and Harstad, 2017; Yaziji and Doh, 2009). Yet, their strategic choices remain poorly understood. Much of the literature treats NGOs as either principled advocates motivated by social good (Besley and Ghatak, 2005; Bénabou and Tirole, 2006; Daubanes and Rochet, 2019) or as dysfunctional entities prone to corporate capture and strategic manipulation (Grossman and Helpman, 1994; Bertrand et al., 2020; Akey et al., 2021). We argue and demonstrate empirically that NGOs adapt their strategies in response to incentives, shaping which stakeholders they engage and ultimately affecting the social and environmental performance of targeted firms.
To study how NGOs respond to incentives and their implications for firms, we focus on a setting where firms are especially exposed to outside pressure: the Annual General Meeting (AGM). AGMs are predictable focal points at which executives present results, shareholders vote on proposals, and media attention peaks (Fioretti et al., 2025). Figure 1 shows that NGO campaigns cluster heavily around AGM weeks: incidence is four to six times higher than in non-AGM weeks (Panel a), and media coverage spikes simultaneously (Panel b). These striking patterns raise a central question: do NGOs target AGMs mainly to maximize exposure, or do they also use campaigns to exert influence over corporate outcomes?
@@
Examining campaign timing reveals that younger NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020). These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents. For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025). Indeed, we find little evidence that other stakeholders are moved by these campaigns. For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves. Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively young. These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times. This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
As NGOs accumulate visibility through AGM-date campaigns, their strategy changes. Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes. By contrast, campaigns in the months preceding the AGM raise the likelihood of proposal success. Campaigns one to three months before the meeting increase support by 8%, while earlier ones often induce settlements that improve firms’ environmental and social performance.
Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns earlier in time as they gain AGM exposure. First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions. This pattern is especially pronounced among NGOs with high search intensity, which may not drive substantial visibility benefits from AGM-day campaigns.
Repeated targeting of AGMs also plays a distinct role. By returning to these focal moments, NGOs build a “reputation for toughness”—demonstrating their ability to amplify negative media coverage when firms are most exposed—which can enhance the effectiveness of subsequent campaigns. Yet, once reputation is established, further AGM-day campaigns yield diminishing benefits, generating smaller increases in NGO search intensity, lower success rates for related shareholder proposals, and less negative media coverage on the target firm. This pattern reflects NGOs’ changing incentives because of evolving stakeholder expectations. As NGOs become recognized and credible actors (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their actions are interpreted differently: what initially signals resolve later signals routine. Established NGOs are therefore expected to engage in more constructive tactics, such as supporting shareholder proposals, rather than relying on visible confrontation.
Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—illustrates how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for their social and environmental impact. In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982), but the meaning and effectiveness of those signals evolve as stakeholders update their expectations about each NGO’s toughness and legitimacy.
This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022; Allcott et al., 2022). More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
Prior studies highlight the role of shareholders influencing firms via voice through voting or exit via divestments (e.g., Cuñat et al., 2016; Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), and of consumers through willingness to pay (Gneezy et al., 2010). We instead show that NGOs act as coordination mechanisms, enabling diverse stakeholders to collectively discipline firms (Koenig and Poncet, 2019; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023). In doing so, NGOs expand the set of channels through which civil society shapes corporate conduct.
Finally, our methodology relates to the literature on the strategic timing of actions to exploit fluctuations in media attention. Prior work shows that politicians and governments time executive orders, military operations, and even disaster relief to coincide with or avoid competing news (Eisensee and Strömberg, 2007; Durante and Zhuravskaya, 2018; Djourelova and Durante, 2022), and that social movements strategically leverage focal dates to amplify visibility (Madestam et al., 2013). We show that NGOs adopt similar tactics in the corporate domain.

### v8

@@ Abstract
This paper studies how the timing of NGO activism shapes its effectiveness in influencing corporate behavior. Using data on 2,500 campaigns targeting U.S. firms, we show that campaigns timed at annual general meetings (AGMs) generate large visibility gains but little contemporaneous influence, while campaigns launched before the AGM significantly increase shareholder proposal success and improve firms’ environmental and social performance. We develop a dynamic model in which NGOs trade off awareness building and credibility formation, generating a lifecycle in activism from visibility-seeking to influence-oriented engagement. Therefore, NGOs’ objectives evolve endogenously to coordinate stakeholder pressure and shape corporate behavior.
@@ 1 Introduction
Non-governmental organizations (NGOs) have become influential actors in shaping firms’ environmental and social policies (e.g., baron2001private; egorov2017private; yaziji2009ngos). Yet, we know remarkably little about their strategic behavior. NGOs are highly opaque: they disclose almost nothing about internal decisions, incentives, or tactics, making most of their actions effectively unobservable (maxwell2012economic). As a result, the literature largely portrays NGOs either as principled advocates driven by social missions (e.g., besley2005competition; benabou2006incentives; daubanes2019rise) or as organizations vulnerable to strategic misuse and capture (e.g., grossman1994protection; bertrand2020tax; akey2021hacking). We argue instead that NGOs respond systematically to incentives generated by stakeholder attention and expectations, shaping NGOs’ strategic choices, with consequences for corporate behavior and social and environmental outcomes.
@@
A rare opportunity to observe these strategic responses arises from campaign timing. NGO campaigns must be public to be effective, and their timing provides one of the few visible choices NGOs make. We focus on how NGOs align campaigns with a moment when firms are uniquely exposed to outside scrutiny: the Annual General Meeting (AGM). AGMs are predictable focal events at which executives report results, shareholders vote, and media attention peaks (fioretti2024shared). Panel (a) of Figure 1 shows that NGO campaigns are five to six times more likely to occur in AGM weeks than at other times, and Panel (b) shows that media coverage of targeted firms peaks at the same moment. These patterns raise a central question: do NGOs target AGMs primarily to maximize exposure, or do they use campaign timing to exert influence on firms? Is there a tradeoff between exposure and influence? More broadly, what incentives shape NGO strategy, and how do their choices affect the firms they target?
Leveraging the AGM as a predictable focal point, we address these questions and uncover a lifecycle in NGO activism. Younger NGOs concentrate on AGM-day campaigns that build visibility, while more established NGOs shift to pre-AGM campaigns that influence shareholder decisions and firms’ environmental and social performance. A simple dynamic model, in which NGOs trade off awareness building against credibility formation, rationalizes this timing pattern. Taken together, our results show that NGO behavior is governed by incentives that determine which stakeholders are mobilized and when, and thus whether campaigns generate symbolic exposure or substantive corporate change. In particular, because most shareholder votes are cast before the AGM, campaigns timed on the meeting date primarily affect visibility, whereas campaigns launched earlier can influence voting outcomes – linking NGO effectiveness to institutional features such as AGM and proxy rules and disclosure timing.
Imagine an NGO uncovers troubling information about a firm and seeks to change its behavior. Because Annual General Meetings (AGMs) occur at predictable times, the NGO can choose whether to campaign before the AGM—when shareholder votes and negotiations are still open—or at the AGM itself, when public and media attention peak but most proxy votes have already been cast. If NGOs simply campaigned as soon as concerns arose, we would not expect the sharp clustering of campaigns around AGMs shown in Figure 1. The pronounced spike instead suggests that campaign timing is a strategic choice.
To study this choice, we assemble a novel dataset of roughly 2,500 campaigns by international NGOs targeting U.S. firms, linked to media coverage, Google search activity, NGO financial statements, shareholder proposals, and environmental and social (E&S) outcomes. SEC rules tie proxy filings and AGM dates to the fiscal year, so each firm’s AGM recurs within a narrow window each year, creating a predictable surge in visibility alongside a fixed deadline for shareholder voting. This institutional structure generates sharp variation in when campaigns occur relative to when shareholder decisions are made, which we exploit to distinguish exposure-oriented from influence-oriented activism and their drivers.
Our first set of results shows that campaign timing is tightly linked to how NGOs affect firms. In their first campaigns, NGOs tend to target the AGM date, generating large visibility shocks – raising media coverage, online search activity, donations, and the likelihood that related shareholder proposals are filed at the next AGM – but having little effect on contemporaneous proposal success or firms’ E&S performance. This pattern suggests that shaming campaigns alone may have limited influence on corporate decisions.
As NGOs gain experience and reputation, campaigns shift into the months before the AGM, when proxy voting and negotiations are still open. These earlier actions significantly increase the success and vote share of E&S proposals and are followed by improvements in firms’ E&S scores, consistent with genuine changes in corporate policy rather than pure reputational damage.
These patterns are not explained by NGOs gradually moving from easier to harder targets: observable characteristics of target firms do not change systematically as NGOs accumulate campaigns. To address selection on unobservables, we exploit quasi-exogenous variation from the publication date of the Fashion Transparency Index, which is fixed by the Rana Plaza anniversary and thus pins down both when campaigns occur and which firms are targeted. Using the distance between index publication and firms’ AGM dates as a source of exogeneous timing variation, we find similar effects both in signs and magnitudes to those estimated in the whole dataset.
We then show that NGOs’ use of AGM visibility versus pre-AGM influence evolves systematically over their organizational life cycle. NGOs with little prior AGM exposure or public visibility concentrate campaigns on the AGM date, where media attention peaks and visibility gains are largest. As NGOs accumulate visibility – proxied by past AGM campaigns, search intensity, and organizational size – they progressively shift campaigns earlier in the AGM cycle, when shareholder votes have not yet been cast and managerial concessions are more likely.
NGO income statement data further indicate that early campaigns are costlier and less visibility-oriented than AGM-day campaigns. For younger NGOs, AGM-day actions generate sizable short-run increases in donations and returns on assets, consistent with shaming campaigns being relatively inexpensive and effective visibility tools. For more established NGOs, the marginal fundraising benefit of AGM campaigns is small, and non-AGM campaigns do not produce comparable financial gains for any group.
The empirical patterns point to a sharp distinction between campaigns launched before the AGM and those timed on the meeting date. Campaigns before the AGM are associated with greater influence on shareholder votes and firm outcomes, while AGM-day campaigns primarily generate visibility. We interpret these findings through a dynamic model in which an NGO, shareholders, and the firm interact repeatedly over time, and in which campaign timing is the NGO’s central strategic choice.
In each period, the public state is summarized by two variables. The first is an awareness stock, which captures how salient the issue is among dispersed shareholders and the broader public. Awareness increases with highly visible actions—most notably campaigns timed on the AGM—and decays over time in the absence of renewed attention. The second is a credibility belief, which reflects stakeholders’ assessment of the NGO’s resolve and effectiveness in sustaining pressure on firms. This belief evolves endogenously as shareholders and firms observe the NGO’s actions and their outcomes.
Given the current state, the NGO chooses whether to campaign before shareholder votes are cast or at the AGM. AGM-day campaigns maximize visibility and generate warm-glow benefits for the NGO, but they occur after most proxy votes are locked in and therefore have little contemporaneous influence on firm decisions. Early campaigns, by contrast, are less visible and forego some warm-glow benefits, but they can affect shareholder voting and negotiations with the firm. Importantly, early campaigns are also more informative: they require sustained engagement and, when successful, reveal persistence and competence, strengthening the NGO’s credibility. AGM-day campaigns are comparatively cheap and visible, but once an NGO is established they convey little new information about its type.
Shareholders observe campaign timing and, when relevant, campaign outcomes. Based on the current levels of awareness and credibility, they decide whether to file a proposal and how much support to give it. Higher awareness increases the likelihood that shareholders notice and coordinate around an issue, while higher credibility raises the expected effectiveness of activism and the probability of success. The firm, anticipating these responses, adjusts its behavior accordingly.
This dynamic interaction yields a simple equilibrium structure. When credibility and awareness are low, the NGO optimally campaigns at the AGM: visibility and warm-glow benefits dominate, and early engagement is unlikely to be persuasive. As awareness accumulates and credibility improves, the NGO crosses a threshold at which the influence and reputational gains from early campaigning outweigh the loss in visibility. At that point, it switches to pre-AGM campaigns to exert direct pressure on shareholder decisions and firm policies. This cutoff behavior reflects evolving stakeholder expectations: actions that initially signal resolve later appear routine, leading established NGOs to rely less on public shaming and more on early, substantive engagement (baur2011moral; mitchell2017reputations).
A purely visibility-driven model would predict persistent AGM-day campaigning. Matching the observed shift from late, high-visibility actions to earlier, influence-oriented engagement instead requires a dynamic reputation-building mechanism in the spirit of intertemporal signaling models (kreps1982reputation; milgrom1982limit; cabral2010dynamics), combined with a deterministic awareness stock that separates visibility-driven exposure from credibility-driven influence. This framework provides a unified explanation for why NGOs initially seek attention and later convert reputation into corporate impact.
Our analysis yields three main contributions. First, we offer a new quantitative framework to examine NGO strategy and to infer how NGOs respond to incentives, which stakeholders they prioritize at different stages, and how their objectives evolve. Classic corporate-finance approaches tie organizational objectives to ownership structure (e.g., grossman1977value; grossman1979theory), while stakeholder theory emphasizes salience based on power, legitimacy, and urgency (freeman1984strategic; mitchell1997toward). Our results show that NGO objectives shift with how donors and investors interpret NGOs’ actions and with credibility concerns about timing. In this sense, we provide a quantitative setting to study how organizational objectives emerge from interactions with multiple stakeholders rather than being fixed by a single principal—for example, an entrepreneur defining the firm’s goals (e.g., scott2002love; romer2006firms; duarte2025competitive; fioretti2024shared). [^1] Our approach also relates to work on the strategic timing of actions around news cycles and focal dates (eisensee2007news; madestam2013do; durante2018attack; djourelova2022media), but shifts the focus to NGOs in the corporate domain and to the management of reputational capital over time.
Second, we contribute to the corporate-finance and stakeholder-governance literatures by identifying mechanisms through which firms partially internalize social and environmental externalities (fioretti2022caring; allcott2022economic). Our setting complements work on shareholder activism and gadflies (e.g., brav2008hedge; gantchev2021costs; malenko2024information): NGOs do not vote but act as external coordinators that put issues on the ballot and help align dispersed investors’ behavior, shaping both the filing and success of shareholder proposals (see also chuah2024shareholder; mazet2023some; mcgahan2023new).
Third, our approach has implications for policy. For boards, early, credibility‐backed campaigns are informative signals that warrant engagement rather than dismissal (dimson2015active; kolbel2017media). For investors and stewardship codes, credible NGOs supply information that helps coordinate dispersed shareholders on E&S issues, complementing traditional voice and exit (oehmke2020theory; gans2021exit; broccardo2020exit; saint2024exit; bonnefon2025moral). For regulators, rules on AGM timing, proposal access, and disclosure shape the returns to NGO campaigns and thus the scope for private activism to substitute for, or complement, formal regulation.